OpenStudy (anonymous):

find the number of horizontal asymptotes of y= 2 - ln x a. 0 b. 1 c. 2 d. 3 e. 4 thankss

OpenStudy (mathteacher1729):

Are you allowed to graph this on a calculator?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

o wait its zero since it is an exponential function that goes to infinity

OpenStudy (agreene):

A. \[\lim_{x\rightarrow\infty}2-\ln(x)=-\infty\] \[\lim_{x\rightarrow-\infty}2-\ln(x)=-\infty\]

OpenStudy (mathteacher1729):

The limit as x goes to negative infinity does not exist (as a real number) because log(x) is not defined (as a real number) for x < 0.

OpenStudy (mathteacher1729):

There is no horizontal asymptote.